OpenStudy (anonymous):

Runners in a long distance race start out going 8 kilometers east and then head north for the remainder of the race. The finish line is 10 kilometers from the starting line. How far did the runners travel north?

OpenStudy (anonymous):

draw a triangle and use pisagor

OpenStudy (asadkarim7):

sqrt 10^2 - 8^2 sqrt 100 - 64 sqrt 36 6 km

OpenStudy (anonymous):

shanks

OpenStudy (anonymous):

you should just draw a triangle and use pisagor theorem, if we say a,b,c edges a=8 and then c = 10 and if the degree between a and b = 90 then to pisagor theorem a^2 + b^2 = c ^2 so that 10^2 = 8^2 + b^2 solve this you will have 6 as the answer ^^ good luck to you